For the first question you must write a short expl
Answer #1

For resistors

It provides resistance to flow of current.

If resistors are grouped in series , than Req = R1 + R2 + R3

If resistors are grouped in parallel , than 1/Req = 1/R1 + 1/R2 + 1/R3

=> V = I * Req

For capacitors

It stores the electric charge .

If capacitors are grouped in series , than 1/Ceq = 1/C1 + 1/C2 + 1/C3

If capacitors are grouped in parallel , than Ceq = C1 + C2 + C3

=> Q = Ceq * V